OpenStudy (anonymous):

On a number line, what is the distance from -10 to -4? Choose one answer. A. 4 units B. 6 units C. 10 units D. 14 units

OpenStudy (acxbox22):

-4-(-10)=-4+10=6 6 units is your answer as it represents th space between -10 and -4

OpenStudy (cwrw238):